# Leverage Cross-Attention for End-to-End Open-Vocabulary Panoptic Reconstruction

### v1

@@ VI-G Scene-level Panoptic Segmentation
In this section, we evaluate the scene panoptic segmentation performance using $PQ$, $SQ$, and $RQ$. $PQ$, a metric that simultaneously reflects both segmentation and recognition quality, is the most indicative of panoptic segmentation performance. PR++ achieves the highest $PQ$ and $SQ$ scores across all sequences and datasets, attributed to the end-to-end learning of panoptic segmentation. As shown in Fig. 6 and Fig. 7, only our method guarantees the consistency between semantic and instance labels, as they are derived by Bayes rule based parameter-free panoptic head. The softmax operation within the panoptic head enables simultaneous consideration of semantic and instance branches during inference and inherently prevents conflicts between the semantic label and instance ID. Only in ”5748ce6f01” scene, PanopticRecon++ underperforms PN. Because PN benefits from the limited overlap ground truth bounding boxes in this sequence.
@@
TABLE IV: Reconstruction quality using different methods on 3 datasets
Acc.$\downarrow$ Com.$\downarrow$ C-L1$\downarrow$ Pre.$\uparrow$ Re.$\uparrow$ F1$\uparrow$
Kimera 1.98 8.39 5.19 88.13 75.06 81.07
@@
The table represents the quality on Replica, ScanNet, and ScanNet++
from top to bottom.
TABLE V: Ablation Studies on ScanNet++ dataset
@@
46.37 73.12 59.81 78.20 88.05 54.30 62.35 73.72 86.83 67.48 65.17
65.73 73.26 82.40 71.61 81.10 56.29 68.71 64.98 75.94 53.36 53.71
76.45 83.41 90.73 79.71 89.90 74.52 78.08 73.02 85.40 84.20 75.47
78.99 84.8 92.37 80.64 90.37 76.01 79.12 75.56 87.31 88.98 79.16
78.99 84.8 92.37 82.80 90.81 76.01 79.12 76.50 87.37 88.98 79.16
VI-H Geometric Reconstruction
To evaluate the quality of scene geometric reconstruction, we select Kimera, PL, and PVLFF as our baseline methods and conduct experiments on three datasets: Replica, ScanNet-V2, and ScanNet++. CL and PR were not chosen as baselines due to their geometric reconstruction approaches being the same as PL and PR++, respectively. Notably, both Replica and ScanNet++ provide depth observations projected from high-precision scene mesh models, while ScanNet-V2 offers noisy depth maps obtained from real depth cameras, and its mesh ground truth is of relatively lower quality. Consequently, the reconstruction quality of all methods on ScanNet-V2 is generally lower compared to the other two datasets. As shown in Tab. IV, our proposed method achieves reconstruction accuracy less than $1cm$ on both Replica and ScanNet++, and less than $3cm$ on ScanNet-V2, significantly outperforming other methods. In contrast to traditional reconstruction methods like Voxblox [59] employed in Kimera, our approach leverages a multi-level feature grid and a compact MLP to represent the SDF surface, enhancing surface smoothness and completeness through regularization terms. While our method places a greater emphasis on surface details, it does not prioritize real-time performance, potentially leading to longer optimization times compared to Voxblox. Both PL and PVLFF are implicit reconstruction methods that rely on density-based representations, which are prone to generating floating spatial artifacts and are less focused on surface quality compared to SDF-based methods. Therefore, considering both panoptic segmentation and reconstruction, the PanopticRecon++ contributes even better performance than the 3D segmentation metrics in Tab. I, Tab. II and Tab. III. Several meshes colorized by panoptic classes and semantic classes are visualized in Fig. 8. As shown in cases, PanopticRecon++ produces smoother wall and floor surfaces, as well as more detailed small objects (e.g. keyboard and mouse).
VI-I Ablation Studies
@@
Segmentation. We analyze the four key design choices of PanopticRecon++: instance spatial prior (Prior), dynamic adjustment for tokens (Adj.), panoptic head (Pan.), and instance-level class feature ($h_{SI}$). The results of ablation studies are shown in Tab. V. By experimenting with a model that includes only the instance spatial prior (Eq. 8), we found that the lack of dynamic adjustment results in over-segmentation, significantly degrading the segmentation performance. This is primarily attributed to the redundancy and overlap of independently optimized binary masks, leading to too many instances. When dynamic adjustment is applied solely, the removal of overlapping and intersecting binary masks results in a notable performance boost. Interestingly, this outcome is similar to the segmentation performance of the PL method. This indicates that relying solely on 2D supervision can result in different objects being mistakenly assigned the same instance ID, especially for objects that do not co-occur in the same image, leading to a substantial decline in segmentation performance. By combining instance spatial prior with dynamic adjustment, our method reveals the advantages of instance tokens and spatial prior, leading to significant improvements in panoptic and instance segmentation performance.
To further validate the effectiveness of the panoptic head in panoptic segmentation and its ability to align semantic and instance labels, we train a complete version of PanopticRecon++ incorporating the panoptic head. We conduct two sets of inference experiments: one queries semantic labels from the semantic branch directly (without $h_{SI}$), and the other replaces instance masks with instance semantic labels (with $h_{SI}$). Our experiments reveal that the choice of inference strategy using instance semantic labels has better semantic segmentation performance. This is due to the avoidance of the negative impact caused by the semantic detection errors of 2D foundation models due to variations in viewpoints. Additionally, both experiments demonstrate an improvement in $PQ$ compared to the model without the panoptic head, which can be attributed to the panoptic head’s capability of Bayes rule based semantic and instance integration.
Rendering. A high-fidelity novel view synthesis (NVS) capability and a color-geometric consistent model are essential for robotic applications. We evaluate NVS performance and the alignment effectiveness between appearance and geometry branches. Alignment quality was assessed using the $AbsRel$ metric on the depth maps rendered by the appearance brunch, while NVS performance was evaluated using $PSNR$, $SSIM$, and $LPIPS$ metrics on the NVS test frames provided in ScanNet++. The results are presented in Fig. 9. Unsurprisingly, the employment of depth improves the model’s NVS capability, which is explained by reducing floating Gaussians in free space. However, the additional task constrains the fitting ability of the appearance model, explaining the minor decrease of $LPIPS$ in novel views.
VI-J User Case
VII Conclusion
We propose PanopticRecon++, an end-to-end open-vocabulary panoptic reconstruction method incorporating multi-branch neural fields and learnable 3D Gaussian-modulated instance tokens. From a cross-attention perspective, we integrate segmentation features with spatial priors, enabling back-propagation of token coordinates and thus addressing the 3D instance ambiguity without post-processing. Explicit instance tokens enable controlled 3D instance initialization and dynamic token adjustment during training. Furthermore, a parameter-free panoptic head ensures consistent and high-precision semantic and instance labels. We evaluate PanopticRecon++’s segmentation, reconstruction, and rendering performance on simulated and real-world datasets. Results demonstrate that our method achieves globally consistent 3D segmentation and reconstruction of 3D scenes, exhibiting competitive performance.

### v2

@@ VI-F Scene-level Semantic Segmentation
PL achieves the highest 2D semantic segmentation scores on KITTI-360 (Tab. III), but only due to superior performance on two broad, ambiguously defined classes, building and vegetation, where its MLP-based smooth boundaries align with the coarse, smoothed ground truth in KITTI-360; for all other classes, our method significantly outperforms it.
@@ VI-G Scene-level Panoptic Segmentation
In this section, we evaluate the scene panoptic segmentation performance using $PQ$, SQSQ, and $RQ$. PQPQ, a metric that simultaneously reflects both segmentation and recognition quality, is the most indicative of panoptic segmentation performance. PR++ achieves the highest $PQ$ and $SQ$ scores across all sequences of indoor datasets, attributed to the end-to-end learning of panoptic segmentation. As shown in Fig. 6 and Fig. 7, only our method guarantees the consistency between semantic and instance labels, as they are derived by Bayes rule based parameter-free panoptic head. The softmax operation within the panoptic head enables simultaneous consideration of semantic and instance branches during inference and inherently prevents conflicts between the semantic label and instance ID. Only in ”5748ce6f01” scene, PanopticRecon++ underperforms PN. PN benefits from the limited overlap ground truth bounding boxes in this sequence. However, PN is provided with the highest-quality 3D polygonal bounding boxes on KITTI-360. As shown in Tab. VI, this privileged access to precise priors enables PN to achieve the highest panoptic segmentation performance. In contrast, PanopticRecon++, operating without any 3D supervision, attains a PQ score competitive with PN.
@@
TABLE IV: Ablation Studies on ScanNet++ dataset
Method 2D-Scene 3D
[…]
46.37 73.12 59.81 78.20 88.05 54.30 62.35 73.72 86.83 67.48 65.17
65.73 73.26 82.40 71.61 81.10 56.29 68.71 64.98 75.94 53.36 53.71
66.80 78.39 85.22 72.84 81.34 64.51 71.88 67.73 85.99 72.46 70.69
74.63 82.42 89.78 79.78 90.21 71.67 76.68 74.21 87.49 84.25 76.51
74.63 82.42 89.78 78.21 89.44 71.67 76.68 70.81 84.58 84.25 76.51
78.99 84.80 92.37 80.64 90.37 76.01 79.12 75.56 87.31 88.98 79.16
78.99 84.80 92.37 82.80 90.81 76.01 79.12 76.50 87.37 88.98 79.16
TABLE V: Reconstruction quality using different methods on 3 datasets
[…]
Replica Kimera 1.98 8.39 5.19 88.13 75.06 81.07
@@
VI-H Parameter Sensitivity
Our parameter settings are fairly robust for instances across most scenes; however, in certain special cases they exhibit some sensitivity and require appropriate adjustment. In practice, we typically adjust the learning rates of the instance tokens, including the coordinate learning rate $l_{xyz}$ and the scaling learning rate $l_{s}$. These parameters determine the movement speed of instance tokens during training as well as their spatial influence. For example, in scene 1ada7a0617 from ScanNet++, using the commonly adopted learning-rate settings (mentioned in Sec. VI-D) results in the keyboard on the table not being matched by any instance token, as shown in Fig. 9. To address this, we set $l_{xyz}=5e{-3}$ and $l_{s}=1e{-2}$ to facilitate faster matching of instance tokens to instances of different scales.
VI-I Geometric Reconstruction
To evaluate the quality of scene geometric reconstruction, we select Kimera, PL, and PVLFF as our baseline methods and conduct experiments on three datasets: Replica, ScanNet-V2, and ScanNet++. CL and PR were not chosen as baselines due to their geometric reconstruction approaches being the same as PL and PR++, respectively. Notably, both Replica and ScanNet++ provide depth observations projected from high-precision scene mesh models, while ScanNet-V2 offers noisy depth maps obtained from real depth cameras, and its mesh ground truth is of relatively lower quality. Consequently, the reconstruction quality of all methods on ScanNet-V2 is generally lower compared to the other two datasets. As shown in Tab. V, our proposed method achieves reconstruction accuracy less than $1cm$ on both Replica and ScanNet++, and less than $3cm$ on ScanNet-V2, significantly outperforming other methods. In contrast to traditional reconstruction methods like Voxblox [65] employed in Kimera, our approach leverages a multi-level feature grid and a compact MLP to represent the SDF surface, enhancing surface smoothness and completeness through regularization terms. While our method places a greater emphasis on surface details, it does not prioritize real-time performance, potentially leading to longer optimization times compared to Voxblox. Both PL and PVLFF are implicit reconstruction methods that rely on density-based representations, which are prone to generating floating spatial artifacts and are less focused on surface quality compared to SDF-based methods. Therefore, considering both panoptic segmentation and reconstruction, the PanopticRecon++ contributes even better performance than the 3D segmentation metrics in Tab. I, Tab. II and Tab. III. Several meshes colorized by panoptic classes and semantic classes are visualized in Fig. 8. As shown in cases, PanopticRecon++ produces smoother wall and floor surfaces, as well as more detailed small objects (e.g. keyboard and mouse).
Figure 9: Comparison of instance token learning rates before and after adjustment.
VI-J Ablation Studies
We analyze the five key design choices of PanopticRecon++: 3D Gaussian-modulated token representation (3DG), instance spatial prior (Prior), dynamic adjustment for tokens (Adj), panoptic head (Pan), and instance-level class feature ($h_{SI}$). The results of ablation studies are shown in Tab. IV.
Token Representation. Table IV shows that 3D Gaussian-modulated tokens (3DG) outperform point cloud-based instance tokens for segmentation (rows 4 vs. 3). This is because point cloud-based tokens, which rely on Euclidean distances, poorly represent objects with significant scale and geometric variation, such as poles, keyboards, and elongated tables. Euclidean distance inadequately captures instances’ spatial extent, orientation, and structure. In contrast, the 3D Gaussian-modulated token explicitly learns the spatial distribution of each instance, naturally accommodating differences in scale and shape. This leads to a more principled and discriminative modeling of inter-instance geometric relationships, resulting in significantly improved segmentation performance.
Spatial Prior. The comparison between the second and fourth rows in Tab. IV clearly demonstrates the critical contribution of spatial prior (Eq. 8) to improved segmentation performance. Interestingly, the result in the second row closely matches the segmentation performance of PL method. This indicates that relying solely on 2D supervision can result in different objects being mistakenly assigned the same instance ID, especially for objects that do not co-occur in the same image, leading to a substantial decline in segmentation performance.
Dynamic Adjustment. As demonstrated by the comparison between the first and fourth rows in Tab. IV, we found that the lack of dynamic adjustment results in over-segmentation, significantly degrading the segmentation performance. This is primarily attributed to the redundancy and overlap of independently optimized binary masks, leading to too many instances. When dynamic adjustment is applied solely, the removal of overlapping and intersecting binary masks results in a notable performance boost. By combining instance spatial prior with dynamic adjustment, our method reveals the advantages of instance tokens and spatial prior, leading to significant improvements in panoptic and instance segmentation performance.
Panoptic Head and Instance-level Class. The spatial prior and dynamic token adjustment are indispensable; removing them leads to a severe performance drop (Tab. IV, rows 1–2). The panoptic head and the instance-level class feature $h_{SI}$, in contrast, provide complementary enhancements. To evaluate their contributions, we fix the prior and adjustment and compare settings with/without the panoptic head and $h_{SI}$. Results indicate that $h_{SI}$ effectively refines noisy semantic predictions when instance segmentation is reliable, but misaligned tokens propagate errors, yielding performance even inferior to supervision from 2D semantics alone. Additionally, both experiments demonstrate an improvement in $PQ$ compared to the model without the panoptic head, which can be attributed to the panoptic head’s capability of Bayes rule based semantic and instance integration.
VI-K User Case
VI-L Failure Case and Limitation
As shown in Fig. 10, the limited generalization of 2D foundation models sometimes restricts our system. On the left, GroundedSAM exhibits typical errors: the same cabinet is labeled inconsistently (“cabinet” / “door”), a door is segmented with varying granularity due to its frame, and the trash bin is misclassified as a toilet. On the right, our method shows robustness to granularity variation (e.g., consistent cabinet instance), but frequent semantic mislabeling still propagates into instance tokens. As a result, although fragmented 2D labels are alleviated, the entire instance may still be assigned an incorrect semantic label.
Another limitation of our method is the high GPU memory consumption due to the large scene containing many objects. This is because the memory cost of our attention mechanism scales with the number of instances.
Figure 10: Failure Case: 3D Segmentation Errors Caused by GroundedSAM Mislabeling
VII Conclusion
We propose PanopticRecon++, an end-to-end open-vocabulary panoptic reconstruction method with multi-branch neural fields and 3D Gaussian-modulated instance tokens. Through cross-attention, segmentation features are integrated with spatial priors, enabling token coordinate back-propagation and resolving 3D instance ambiguity without post-processing. Learnable tokens allow controlled initialization and dynamic adjustment, while a parameter-free panoptic head ensures consistent semantic and instance labels. Experiments on simulated and real datasets show that PanopticRecon++ achieves globally consistent 3D segmentation and reconstruction with competitive performance. In future work, we plan to distill the entire scene into 3DGS to achieve real-time performance for robotic systems.
Appendix
VII-A Generalization to Large-scale Scenes
TABLE VI: Panoptic Segmentation quality using different methods on KITTI-360 dataset
Method 2D-Scene 3D
@@ VII-A Generalization to Large-scale Scenes
Close-vocabulary:
Panoptic NeRF 62.26 68.97 78.16 64.39 74.59 74.99 91.96 61.24 69.91 88.05 92.30
Open-vocabulary:
Panoptic Lifting 42.50 54.79 55.33 58.06 74.20 28.86 41.99 54.16 72.93 21.93 29.44
PVLFF 45.20 52.76 57.69 54.81 61.69 13.33 23.89 26.69 36.65 13.59 15.87
PanopticRecon 57.61 66.51 74.16 56.21 69.86 64.42 85.98 62.60 80.59 86.36 85.99
PanopticRecon++ 60.23 68.34 74.51 57.36 71.66 68.59 86.39 67.89 82.42 91.26 87.46
Figure 11: Comparison of the quality of semantic segmentation and panoptic meshes of different methods on KITTI-360.
Figure 12: Comparison of the quality of semantic segmentation and panoptic meshes of different methods on KITTI-360.
To validate the generalization ability of our method across varying scene scales and sensor modalities, we extend our evaluation to outdoor large-scale scenes, utilizing RGB images and LiDAR point clouds as joint multimodal inputs. Specifically, we select a 200m segment of sequence 00 from the KITTI-360 dataset for this evaluation. The sensor configuration includes depth observations from a Velodyne HDL-64 LiDAR and a single perspective camera. In our experiments, the set of semantic labels comprises: road, sidewalk, building, fence, vegetation, car, bicycle, and motorcycle. Notably, car and motorcycle are designated as instance-level categories.
As shown in Tab. VI, Panoptic NeRF, as the closed-vocabulary approach, achieves the highest panoptic segmentation performance, as it is trained using known 3D bounding boxes and precisely aligned semantic classes, effectively operating under the assumption of perfect instance ground truth. Consequently, its results serve as a strong upper bound on achievable segmentation accuracy. Our open-vocabulary method achieves a PQ of 60.23%, just 2% shy of Panoptic NeRF’s performance.
Crucially, the primary advantage of open-vocabulary methods lies in their ability to segment objects absent from the training set and to produce finer, more accurate segmentation boundaries. For instance, as shown in Fig. 13, the bicycle was omitted from the ground-truth annotations; however, the 2D VLM successfully detected it, enabling our open-vocabulary framework to recover its shape and spatial extent accurately.
Figure 13: Comparison of the quality of semantic segmentation and panoptic meshes of different methods on KITTI-360.
Among panoptic segmentation metrics, our method achieves the highest PQ, while Panoptic Lifting attains the highest semantic segmentation metrics. This discrepancy arises because, among all semantic classes, only building and vegetation achieve higher segmentation performance than our method. In particular, vegetation shows a significant performance advantage over ours. For all other classes, Panoptic Lifting performs worse than our approach. This behavior stems from Panoptic Lifting’s reliance on an MLP-based representation, which produces smoother, more regularized boundaries. Furthermore, the ground-truth annotations in KITTI-360 are themselves smoothed and less precise than the true object boundaries in the raw observations. Consequently, Panoptic Lifting benefits from this alignment between its inherent smoothness and the annotated ground truth, particularly for large, ambiguously bounded regions such as vegetation and building, leading to superior performance on these two categories.
VII-B Impact of 2D VLM Errors
Evaluate on unseen categories: We introduce an additional ablation study to evaluate the influence of different text prompt lists on the segmentation performance of our method. Specifically, we conduct experiments on the ScanNet scene 0087_02 using three distinct text prompt configurations: (a) a baseline prompt list augmented with plant; (b) an extended list including plant and box; (c) an extended list further including plant, box, and shelf; and (d) the original prompt list from our manuscript, serving as the baseline. This comparison enables a systematic analysis of how vocabulary expansion affects open-vocabulary segmentation accuracy and robustness.
TABLE VII: Evaluate on Unseen Categories with Different Text Prompt Lists
Method 2D-Scene 3D
@@ VII-B Impact of 2D VLM Errors
a 81.84 88.71 92.38 81.49 87.96 84.81 90.54 80.27 87.28 98.24 86.65
b 79.64 88.34 90.15 81.02 86.59 84.21 89.48 78.50 82.88 94.63 86.34
c 77.52 77.52 87.50 74.39 79.51 83.20 89.33 71.41 80.39 89.28 85.30
d 81.84 88.71 92.38 81.49 87.96 84.81 90.54 80.27 87.28 98.24 86.65
Since the scene 0087_02 contains no instances of plant nor any semantically similar objects, the addition of plant to the text prompt list had no measurable impact on the 2D VLM’s mask predictions, resulting in identical segmentation performance between (a) and (d) as shown in Tab VII. In contrast, (b) introduces an additional prompt box, which triggers only minor false positives in a small subset of images where partial views of tables are observed, resulting in a marginal degradation of segmentation metrics. (c) further incorporates the shelf prompt in addition to (b), increasing misclassifications: in numerous images where doors are partially occluded or incompletely observed, the 2D vision-language model (VLM) erroneously assigns shelf labels due to structural similarities between door frames and shelf geometries. This proliferation of false positives compromises the accuracy of the segmentation output, resulting in a drop in overall metrics compared to (b). These results demonstrate that while text prompts for 2D VLMs should be carefully curated to maximize segmentation fidelity, the system remains robust even under suboptimal prompt selection.
TABLE VIII: Evaluate on Unseen Categories with Different Text Prompt Lists
[…]
Panoptic Lifting 35M 14.5s 15G
PVLFF 110M 12s 10G
Panoptic Recon++ 900M 4s 30G
VII-C Computational Efficiency
We evaluate the model size, per-frame inference time, and memory consumption of our method and competing approaches directly on KITTI-360 using an NVIDIA A6000 GPU. All methods are tested on images of resolution 376×1408. As shown in Tab. VIII, Panoptic Lifting is based on MLP representations, which offer compact model sizes but suffer from slow inference. Both PVLFF and Panoptic Recon++ employ implicit representations over grid-based structures; however, PVLFF incurs additional inference overhead due to costly feature clustering operations. In contrast, our method achieves a substantial acceleration in inference speed, at the cost of a moderate increase in model size.
@@ VII-C Computational Efficiency
VII-D Ablation Study on Rendering
A high-fidelity novel view synthesis (NVS) capability and a color-geometric consistent model are essential for robotic applications. We evaluate NVS performance and the alignment effectiveness between appearance and geometry branches. Alignment quality was assessed using the $AbsRel$ metric on the depth maps rendered by the appearance brunch, while NVS performance was evaluated using $PSNR$, SSIMSSIM, and $LPIPS$ metrics on the NVS test frames provided in ScanNet++. The results are presented in Fig. 14. Unsurprisingly, the employment of depth improves the model’s NVS capability, which is explained by reducing floating Gaussians in free space. However, the additional task constrains the fitting ability of the appearance model, explaining the minor decrease of $LPIPS$ in novel views. We also explore an RGB-D point cloud initialization for 2DGS [66] representation, which achieves promising NVS results as shown in Fig. 15.